OpenStudy (anonymous):

Please help! A car is traveling at a rate of 45 meters per second. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 4 hours? Do not round the answers.

OpenStudy (nikato):

do you know many kilometers in a meter? and how many seconds in an hour?

OpenStudy (anonymous):

1h= 360s 360*45=16200/1000=16.2

OpenStudy (carson889):

Here is a neat little trick to learn. Going from m/s to km/h you multiply the number by 3.6. Going from km/h to m/s you divide the number by 3.6. The reason why would follow from a long hand version of converting 1 km/h or m/s to its converse form. Ex: 1 km/h There are 3600s in one hour (60s/min * 60 min/hour). There are 1000 meters per kilometer. So lets change this km/s from km/h (1 km/h) * (1h/3600s) = (1/3600)km/s Now lets convert to meters: (1/3600) km/s * (1000m / 1 km) = (1000/3600) m/s = (1/3.6) m/s Or: 1 km/h = (1/3.6) m/s To find what 1 m/s equals, simply divide both sides by (1/3.6): (1 km/h) / (1/3.6) = (1/3.6)/(1/3.6) m/s 3.6 km/h = 1 m/s Tada